A liquid causes a pressure of 300000 Pa on an object that is submerged at a depth of 30 m in the liquid? 

 

What is the liquid? 



1
Expert's answer
2021-02-17T10:39:55-0500

The pressure of a liquid at a depth h is:


"p=\rho gh,"


where "\rho" - the density of a liquid;

g - the gravitational acceleration.


To find out what is the liquid, we need to find its density:


"\rho=\frac{p}{gh}=\frac{300000}{9.8\cdot 30} \approx1000\space kg/m^3."


Answer: the liquid is water.